OpenStudy (elisuzsmith):

A box has a volume of x^3 -16x cm3. Another box has a volume of x^3 - 4x^2 cm3. What is the ratio of the volume of the first box to that of the second?

OpenStudy (anonymous):

(x^3-16x)/(x^3-4x^2) =x(x^2-16)/x^2(x-4) =(x+4)(x-4)/x(x-4) =(x+4)/x

OpenStudy (elisuzsmith):

That is what I got, but you don't then cube the answer to find the ratio do you? It just stands the way it is, correct?

OpenStudy (anonymous):

that is the final answer of your question.

OpenStudy (elisuzsmith):

Thank you!

OpenStudy (anonymous):

yw